OpenStudy (phagey):

Subtracting a positive integer is the same as adding a negative integer. True False

OpenStudy (igigighjkl):

Wat grade is tis?

OpenStudy (phagey):

7th

OpenStudy (phagey):

i forgot what the answer was

OpenStudy (zkoqmosii):

Literally tagging people is pointless. The answer is true, but to explain it, you really should get a visual, so I'd check khan academy.

OpenStudy (igigighjkl):

You do khan academy @!!!

TheSmartOne (thesmartone):

Let's say that a and b are both positive numbers a and b a - b = a + (-b) is this true? let's find out so, let's make a = 3, and b = 4 3 - 4 = 3 + (-4) is it true?

TheSmartOne (thesmartone):

hint: x + (-y) = x - y

OpenStudy (igigighjkl):

Sorry I tagged you ALL. /=